OpenStudy (taylor0402):

What is the solution to the system of equations? 5x+4y=8 x-2y=10

OpenStudy (cresciez):

Do you have answer choices we can build off of? c:

OpenStudy (taylor0402):

no sadly

OpenStudy (cresciez):

Alright no biggie :) Do you have any idea what the answers may be?

OpenStudy (cresciez):

If not then let's move on :P

OpenStudy (taylor0402):

ok yes lets move on because i have no idea

OpenStudy (alekos):

Let me help you if you like?

OpenStudy (anonymous):

Okay I love these!

OpenStudy (taylor0402):

im also kind of in a hurry so...yea lets get this done quick

OpenStudy (anonymous):

Okay so lets do this!

OpenStudy (taylor0402):

please =)

OpenStudy (cresciez):

I got it but could you help make sure my explanations are correct? O.o

OpenStudy (anonymous):

Sure haha.

OpenStudy (anonymous):

These are the best :)

OpenStudy (cresciez):

Alright, so 5x+4y=8 x-2y=10 The first step is to multiply everything by 2 except for the end numbers (the equal sign will be the 'boundary') 5x+4y=8 2x-4y=10

OpenStudy (anonymous):

No sorry..

OpenStudy (anonymous):

Thats wrong..

OpenStudy (taylor0402):

oh ok then

OpenStudy (cresciez):

It's not wrong

OpenStudy (anonymous):

The first step is to subtract 2y from each side on the second equation.

OpenStudy (anonymous):

Its much easier that way.

OpenStudy (anonymous):

Then you multiply: When you do that you should get: \[x = 2(y + 5)\]

OpenStudy (anonymous):

so now we know what x is..

OpenStudy (taylor0402):

one sec

OpenStudy (anonymous):

Then you have to plug in x into the first equation... so we would have: \[5(2y +10) + 4y = 8 \]

OpenStudy (anonymous):

When you solve you should get: \[y = -3\]

OpenStudy (cresciez):

If we add my equation we get 7x=28 and x=4. 4-2y = 10 -2y = 10 - 4 -2y = 6 = -3 y = 3 x = 4

OpenStudy (taylor0402):

ok now how do i find x

OpenStudy (anonymous):

Now plug in y into this equation: \[x = 2(y + 5)\]

OpenStudy (taylor0402):

wait so is it 3 or -3

OpenStudy (cresciez):

y = -3 sorry, I forgot that

OpenStudy (anonymous):

So you would solve this to get x: \[x = 2(-3 + 5)\]

OpenStudy (taylor0402):

oh ok

OpenStudy (taylor0402):

oh ok which would be...give me a sec

OpenStudy (anonymous):

Now when you do that you should get: \[x = 4\]

OpenStudy (anonymous):

so the answer is: \[(4, -3)\]

OpenStudy (cresciez):

I hope you get the solution either way Taylor, have a wonderful day ~

OpenStudy (cresciez):

I already said that ...

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @Taylor0402

OpenStudy (cresciez):

???????

OpenStudy (anonymous):

Yes, but you got positive 3 for y... @Cresciez

OpenStudy (anonymous):

Great try though!

OpenStudy (anonymous):

:)

OpenStudy (taylor0402):

thank you so much

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

Have an amazing day! @Taylor0402

OpenStudy (taylor0402):

wow both of you guys did a great job

OpenStudy (taylor0402):

you to @Tom_Boy_Rebel